### v1

@@ 2 Paley ETFs
$\left\{\,f\in\operatorname{AGL}(\mathbb{F}_{p})\,\,\middle|\,\,f(x)=mx+b,\,\,a\in\operatorname{QR}(\mathbb{F}_{p}),\,\,b\in\mathbb{F}_{p}\,\right\}.$
By Proposition 3.1 in [Kan69], since $\{M^{k}T^{\ell}\}$ is isomorphic to the given subgroup of $\operatorname{AGL}(\mathbb{F}_{q})$, it suffices to show that it is the automorphism group and that it acts transitively as the automorphism group in order to show that it acts double-homogeneously but not double-transitively on either the vectors or the lines.
Recall that we denote the automorphism group of the vectors as $G$ and the automorphism group of the lines as $[G]$. Since we will implicitly show that the line automorphisms are $G\times U(1)$, we will treat elements of $[G]$ as being in $\operatorname{U}((p-1)/2)$ rather than $\operatorname{PU}((p-1)/2)$.
Note that Proposition 2.5 implies $G$ and $[G]$ are transitive since the vectors in $\Phi_{p}$ are the orbit of $\mathbbm{1}$ under $\{M^{k}\}$. We now show that $G$ contains $\{M^{k}T^{\ell}\}$. Trivially, the subgroup $\{T^{\ell}\}$ fixes $\mathbbm{1}$. Thus, for an arbitrary vector $M^{\kappa}\mathbbm{1}$ in $\Phi_{p}$, we have
$M^{k}T^{\ell}M^{\kappa}\mathbbm{1}=M^{k}M^{\kappa a^{-\ell}}T^{\ell}\mathbbm{1}=M^{k+\kappa a^{-\ell}}\mathbbm{1}\in\Phi_{p}$
From this it follows that $[G]\supseteq G\supseteq\{M^{k}T^{\ell}\}$.
Now we want to show that there is nothing else in $[G]$. The stabilizer of $\mathbbm{1}\mathbbm{1}^{\top}$ in $\{M^{k}T^{\ell}\}$ is just $\{T^{\ell}\}$. Since $[G]$ is transitive, by the orbit stabilizer theorem, the index of the stabilizer of $\mathbbm{1}\mathbbm{1}^{\top}$ in $[G]$, denoted $[G]_{\mathbbm{1}\mathbbm{1}^{\top}}$, is $p$. The index of $\{T^{\ell}\}$ in $\{M^{k}T^{\ell}\}$ is $p$. Thus, $[G]$ strictly contains $\{M^{k}T^{\ell}\}$ if and only if the stabilizer group of $\mathbbm{1}\mathbbm{1}^{\top}$ strictly contains $\{T^{\ell}\}$. Let $U\in[G]_{\mathbbm{1}\mathbbm{1}^{\top}}$. We seek to show that it is $T^{\ell}$ for some $\ell$ and thus $[G]=\{M^{k}T^{\ell}\}$, which would then imply that $G=\{M^{k}T^{\ell}\}$. Let $\sigma\in S_{0,1,\ldots,p-1}$ be the permutation of the $\varphi_{j}\varphi_{j}^{\ast}$ induced by $U$ (which necessarily fixes 00). By Theorem 1.14, we know that such a $\sigma$ must map triple products with $\operatorname*{TP}(0,k,\ell)$ to those $\operatorname*{TP}(0,\sigma(k),\sigma(\ell))$ of the same value. Plugging $j=0$ into Lemma 2.6, we have
$\operatorname*{TP}(0,k,\ell)=\left\{\begin{array}[]{lr}\alpha^{3}&k\in\operatorname{NR},\ell\in\operatorname{QR},k-\ell\in\operatorname{QR}\\[2.0pt] \left|\alpha\right|^{2}\alpha&(k,k-\ell\in\operatorname{NR},\,\ell\in\operatorname{QR})\textrm{ or }(k,\ell\in\operatorname{NR},\,k-\ell\in\operatorname{QR})\textrm{ or }(k,\ell,k-\ell\in\operatorname{QR})\\[2.0pt] \left|\alpha\right|^{2}\overline{\alpha}&(k,k-\ell\in\operatorname{QR},\,\ell\in\operatorname{NR})\textrm{ or }(k,\ell\in\operatorname{QR},\,k-\ell\in\operatorname{NR})\textrm{ or }(k,\ell,k-\ell\in\operatorname{NR})\\[2.0pt] \overline{\alpha}^{3}&\ k\in\operatorname{QR},\ell\in\operatorname{NR},k-\ell\in\operatorname{NR}\end{array}\right.$
For case $\alpha^{3}$, for any choice of $k\in\operatorname{NR}$, there is always at least one $\ell\in\operatorname{QR}$ such that $k-\ell\in\operatorname{QR}$ and vice versa. Thus, $\sigma$ must map $\operatorname{QR}$ to $\operatorname{QR}$ and $\operatorname{NR}$ to $\operatorname{NR}$. Now consider arbitrary $k\in\operatorname{QR}$ and $\ell=-k\in\operatorname{NR}$. The triple product is $\left|\alpha\right|^{2}\overline{\alpha}$ precisely when $k-\ell=2k\in\operatorname{QR}$, i.e., when $2\in\operatorname{QR}$. This happens when $p\equiv-1\bmod 8$. Otherwise, when $p\equiv 3\bmod 8$, the triple product is $\overline{\alpha}^{3}$. Either way, once the ordering of $\operatorname{NR}$ is fixed to be the ordering of negative the $\operatorname{QR}$, $\sigma$ induces the same permutation of the two sets (of order at most $(p-1)/2$). To that end, let $\Phi_{\operatorname{QR}}$ be the $(p-1)/2\times(p-1)/2$ circulant matrix that has the $\operatorname{QR}$ columns in order
$M^{1}\mathbbm{1},\,M^{(p-3)/2}\mathbbm{1},\,M^{(p-5)/2}\mathbbm{1},\,\ldots,\,M^{3}\mathbbm{1},\,M^{2}\mathbbm{1}.$
Define $\Phi_{\operatorname{NR}}=\overline{\Phi_{\operatorname{QR}}}$, the component-wise complex conjugation. Then $\Phi_{\operatorname{NR}}$ is also circulant and each column corresponds to the negative of the parameter of the column in $\Phi_{\operatorname{QR}}$. By Lemma 2.7, $\Phi_{\operatorname{QR}}$ and $\Phi_{\operatorname{NR}}$ are invertible and their inverses have constant row sums. Let $\Pi$ denote the $(p-1)/2\times(p-1)/2$ permutation matrix representing the permutation of the quadratic residues (respectively, the non residues). We would like to show that $U$ is this permutation (up to unimodular constant). On the vector level, in order for $U$ to induce the given permutations, there must exist diagonal unitary matrices $D_{\operatorname{QR}}$ and $D_{\operatorname{NR}}$ such that
$U=\Phi_{\operatorname{QR}}D_{\operatorname{QR}}\Pi\Phi_{\operatorname{QR}}^{-1}=\Phi_{\operatorname{NR}}D_{\operatorname{NR}}\Pi\Phi_{\operatorname{NR}}^{-1}$
Since $U\mathbbm{1}\mathbbm{1}^{\top}U^{\ast}=\mathbbm{1}\mathbbm{1}^{\top}$, $D_{\operatorname{QR}}$ and $D_{\operatorname{NR}}$ must be unimodular multiples of the identity and can be ignored. Thus, we have
$\left(\Phi_{\operatorname{NR}}^{-1}\Phi_{\operatorname{QR}}\right)\Pi=\Pi\left(\Phi_{\operatorname{NR}}^{-1}\Phi_{\operatorname{QR}}\right).$
As product of circulant matrices, $\Phi_{\operatorname{NR}}^{-1}\Phi_{\operatorname{QR}}$ is circulant and diagonalizable by the Fourier matrix for $\mathbb{Z}_{(p-1)/3}$. Since $\Pi$ commutes with it, it must also be diagonalizable by the Fourier matrix and hence circulant. The only circulant permutation matrices are of the form $T^{\ell}$ and the product commutes:
$U=\Phi_{\operatorname{QR}}T^{\ell}\Phi_{\operatorname{QR}}^{-1}=\Phi_{\operatorname{QR}}\Phi_{\operatorname{QR}}^{-1}T^{\ell}=T^{\ell},$
as desired. ∎
This theorem should not be incredibly surprising, given that Paley designs are the only $2$-homogeneous but not $2$-transitive Hadamard designs [Kan69]. As we will see in Theorem 3.6, Paley skew-conference matrices also yield an infinite class of ETFs which are $3$-homogeneous but not $3$-transitive. In [Kan69], it is also noted that if a set has a $2$-homogeneous automorphism group that is not $2$-transitive, represented as a subgroup of $\operatorname{AGL}(\mathbb{F}_{p})$, then the stabilizer of 00 (for us $\mathbbm{1}$ or $\mathbbm{1}\mathbbm{1}^{\top})$ permutes elements corresponding to the quadratic residues and elements corresponding the non-quadratic residues, which is what we observed.
@@ 3 $k$-Homogeneous ETFs
In this section, we characterize $k$-homogeneous ETFs. Throughout this section, we assume $d>1$. In particular, this means that the ETFs cannot trivially contain repeated elements.
@@
The line automorphism group of orthonormal bases and simplex ETFs with $n$ vectors is isomorphic to $S_{n}$. If an ETF is a regular simplex ETF or an orthonormal basis, then the vector automorphism group is also isomorphic to $S_{n}$.
@@
Given any two ordered orthonormal bases $U_{1}$ and $U_{2}$ of $\mathbb{C}^{n}$, $U_{2}U_{1}^{-1}$ is a unitary matrix that maps the basis $U_{1}$ to the basis $U_{2}$. Thus, the automorphism group of the vectors and in turn the lines is $S_{n}$. A $(d,n)$-ETF has $n=d+1$ vectors if and only if all of the triple products are negative [FJKM18]. Thus, by Theorem 1.14, any permutation of the lines is induced by some element of the automorphism group of the lines. Finally, if the ETF is a regular simplex ETF, then the inner products are all negative and thus any permutation of the vectors is possible (see, e.g., [CW16]). ∎
@@
The only ETFs with $3$-transitive line automorphism groups are orthonormal bases and simplices. The only ETFs with $2$-transitive vector automorphism groups are orthonormal bases and regular simplex ETFs. Finally, the only real ETFs with $2$-homogeneous vector automorphism groups with are orthonormal bases and regular simplex ETFs.
@@
The first result is Theorem II.6 in [Kin19]. The second comes from the fact that if the vector automorphism group is $2$-transitive, then any inner product of distinct vectors is mapped to any other inner product. The third comes from the fact that for real vectors the inner product is symmetric, meaning that $2$-homogeneity implies that all of the inner products between distinct vectors are the same. ∎
Note that this gives an alternate proof that the vector automorphism group of the prime order Paley ETFs $\Phi_{p}$ is not $2$-transitive. We note that the following immediately follows from the definition of homogeneity, but we record it here as we will use it frequently.
@@
If a group acts $k$-homogeneously on a set of size $n$, then it also acts $(n-k)$-homogeneously.
Contrast this with transitivity, where every $k$-transitive group action is also $(k-1)$-transitive and where the larger $k$ is, the less likely it is that a group acts $k$-transitively. We now have a final lemma before stating and proving the main result of this section.
@@
Assume that $\Phi$ is an ETF with $n>3$ vectors where the inner products of any two distinct vectors is imaginary. Then the line automorphism group of $\Phi$ is $3$-homogeneous. Conversely, if $n\notin\{d,d+1,2d\}$, $\Phi$ cannot have $3$-homogeneous automorphism group.
@@
Let $j,k,\ell$ be distinct indices. Then note that
@@
If the inner products are all imaginary, for any set $\{j,k,\ell\}$ of distinct indices there is some ordering of the triple products which is a positive real number times $i$. It now follows from Theorem 1.14 that the automorphism group of the lines is $3$-homogeneous.
Similarly, we note that for any ETF and any set $\{j,k,\ell\}$ of distinct indices, there are only two possible triple products regardless of order, $\operatorname*{TP}(j,k,\ell)$ and $\overline{\operatorname*{TP}(j,k,\ell)}$. Thus, if the triple products are not all complex conjugates of each other, there will be two sets of three distinct indices which cannot be mapped to each other. This means that the triple products of a $3$-homogeneous frame are either real or imaginary. Note that if the triple products are all positive then the vectors are switching equivalent to a set with all positive inner products and thus have an invertible Gram matrix. But the only non-redundant ETFs are orthonormal bases, a contradiction. Thus, the triple products are all $-\alpha$ or all ±$\pm i\alpha$ for some $\alpha\geq 0$. A necessary condition for an ETF to have purely imaginary inner products (see, e.g., [ET00, Str08]) is that $n=2d$. Simplices and orthonormal bases uniquely yield triple products which are all negative or 00, respectively. ∎
In [BP05, HS12], it was proven that redundant equiangular tight frames which are in a certain sense optimally robust to three erasures (called $3$-uniform) must be a simplex or have purely imaginary inner products, which suggests a connection to being $3$-homogeneous.
In [IM22], it was proven that any redundant spanning set which has a doubly transitive line automorphism group must be an ETF. However, the result holds for doubly homogeneous ETFs with minor changes to the proof.
Proposition 3.5.
Let $\Phi=(\varphi_{j})_{j=1}^{n}\in\mathbb{F}^{d\times n}$ be a frame (i.e., spanning set) with $n>\max\{2,d\}$. If $\Phi$ has a doubly homogeneous line automorphism group, then $\Phi$ is an ETF.

### v2

@@ 2.4 Paley ETFs over Fields of Prime Power Order
$\left(\left(\mathbb{Z}_{p}\right)^{s}\rtimes\mathbb{Z}_{(q-1)/2}\right)\rtimes\mathbb{Z}_{s}\cong\left\{\,f\in\operatorname{A\Gamma L}(\mathbb{F}_{q})\,\,\middle|\,\,f(x)=mx^{\sigma}+b,\,\,b\in\mathbb{F}_{q},\,\,m\in\operatorname{QR}(\mathbb{F}_{q}),\,\,\sigma\in\operatorname{Aut}(\mathbb{F}_{q})\,\right\}.$
Thus, $G\leq\operatorname*{Sym}_{v}(\Phi_{q})\leq\operatorname*{Sym}_{\ell}(\Phi_{q})$.
The remainder of the proof is almost exactly like the end of the proof of Theorem 2.10. Namely, consider $x\neq y$ and $z\neq w$ in $\mathbb{F}_{q}$. Then either $S^{-1}(x-y)$ or $S^{-1}(y-x)$ is a quadratic residue, similarly for $(z,w)$. Thus, without loss of generality, we pick an ordering so that $S^{-1}(x-y),S^{-1}(z-w)\in\operatorname{QR}$. Then $M^{\{-y\}}\varphi_{y}=M^{\{-w\}}\varphi_{w}=\mathbbm{1}$, and $M^{\{-y\}}\varphi_{x}$ and $M^{\{-w\}}\varphi_{z}$ are both columns indexed by $S(\operatorname{QR})$. We can then apply a translation $T^{\ell}$ as needed, which fixes $\mathbbm{1}$ and maps $M^{\{-w\}}\varphi_{z}$ to $M^{\{-y\}}\varphi_{x}$. Thus, $M^{\{y\}}T^{-\ell}M^{-\{w\}}$ maps $\{\varphi_{z},\varphi_{w}\}$ to $\{\varphi_{x},\varphi_{y}\}$. Hence, $\operatorname*{Sym}_{v}(\Phi_{q})\leq\operatorname*{Sym}_{\ell}(\Phi_{q})$ act $2$-homogeneously on the vectors and lines, respectively.
We know from [IM24] that $\operatorname*{Sym}_{\ell}(\Phi_{q})$ is not $2$-transitive and thus neither is $\operatorname*{Sym}_{v}(\Phi_{q})$. It thus follows from Lemma 2.9 that $\operatorname*{Sym}_{v}(\Phi_{q})=\operatorname*{Sym}_{\ell}(\Phi_{q})\cong G$. ∎
Corollary 2.20.
Let $q>3$ be a prime power with $q\equiv 3\bmod 4$. The bender of the Paley ETF $\Phi_{q}$ is a BIBD.
Proof.
This follows immediately from Theorem 2.19 and Proposition 1.23
Due to a result from Chebotarëv [SL96], the Paley ETFs $\Phi_{p}$ with $p$ prime are full spark, i.e., have spark (p1)/2+1=(p+1)/2(p-1)/2+1=(p+1)/2. Their bender is just trivially the set of all subsets of size $(p+1)/2$, which is a $t$-design for all $1\leq t\leq(p+1)/2$. For a discussion of $p^{s}$ with $s>1$, see Question 4.1.
@@ 3 $k$-Homogeneous ETFs
In this section, we characterize $k$-homogeneous ETFs. Throughout this section, we assume $d>1$ as otherwise $d>1$ would need to be added to the statement of almost every result. In particular, this means that the ETFs cannot trivially contain repeated elements.
@@
The line symmetry group of orthonormal bases and simplex ETFs with $n$ vectors is $S_{n}$. If an ETF is a regular simplex ETF or an orthonormal basis, then the vector symmetry group is also isomorphic to $S_{n}$.
@@
Given any two ordered orthonormal bases $U_{1}$ and $U_{2}$ of $\mathbb{C}^{n}$, $U_{2}U_{1}^{-1}$ is a unitary matrix that maps the basis $U_{1}$ to the basis $U_{2}$. Thus, the automorphism group of the vectors and in turn the lines is $S_{n}$. A $(d,n)$-ETF has $n=d+1$ vectors if and only if all of the triple products are negative [FJKM18]. Thus, by Theorem 1.14, any permutation of the lines is in the symmetry group. Finally, if the ETF is a regular simplex ETF, then the inner products are all negative and thus any permutation of the vectors is possible (see, e.g., [CW16]). ∎
@@
The only ETFs with $3$-transitive line symmetry groups are orthonormal bases and simplices. The only ETFs with $2$-transitive vector symmetry groups are orthonormal bases and regular simplex ETFs.
@@
The first result is Theorem II.6 in [Kin19]. The second comes from the fact that if the vector symmetry group is $2$-transitive, then any inner product of distinct vectors is mapped to any other inner product. ∎
Note that this gives an alternate proof that the vector symmetry group of the prime order Paley ETFs $\Phi_{p}$ is not $2$-transitive. We note that the following immediately follows from the definition of homogeneity, but we record it here as we will use it frequently.
@@
If a group acts $k$-homogeneously on a set of size $n$, then it also acts (nk)(n-k)-homogeneously.
We now have a final lemma before stating and proving the main result of this section.
@@
If $\Phi$ is an ETF with $n>3$ vectors which has $3$-homogeneous line symmetry group, then up to switching equivalence, the inner products are all 0 ($n=d$), negative ($n=d+1$), or purely imaginary ($n=2d$).
@@
Let $\Phi=(\varphi_{j})_{j}$ be arbitrary and let $j,k,\ell$ be indices. Note that
@@
Also note that triple products are invariant to cyclic permutation of the indices; i.e.,
$\operatorname*{TP}(j,k,\ell)=\operatorname*{TP}(\ell,j,k)=\operatorname*{TP}(k,\ell,j).$
Thus, for any fixed set of indices $\{j,k,\ell\}$, there are only two possible triple products regardless of order, $\operatorname*{TP}(j,k,\ell)$ and $\overline{\operatorname*{TP}(j,k,\ell)}$. Now assume that $\Phi$ is an ETF with $n>3$ vectors which has a $3$-homogeneous line symmetry group. Since $\Phi$ is an ETF, every triple product (of distinct indices) has modulus $\left|a\right|^{3}$ for some $a\in\mathbb{C}$. And further, since $\Phi$ is $3$-homogeneous and every set of three lines is mapped to every other set of three lines, $\operatorname*{TP}(j,k,\ell)\in\{\operatorname*{TP}(\tilde{j},\tilde{k},\tilde{\ell}),\,\overline{\operatorname*{TP}(\tilde{j},\tilde{k},\tilde{\ell})}\}$ for any distinct triples of distinct indices. That is, every triple product must be of the form $a^{3}$ or $\overline{a^{3}}$. Assume without loss of generality that the norms of the vectors of the ETF are all $1$. An ETF with unit norm vectors such that the spectral norm of
$\begin{pmatrix}1&\left\langle\varphi_{j},\varphi_{k}\right\rangle&\left\langle\varphi_{j},\varphi_{\ell}\right\rangle\\ \left\langle\varphi_{k},\varphi_{j}\right\rangle&1&\left\langle\varphi_{k},\varphi_{\ell}\right\rangle\\ \left\langle\varphi_{\ell},\varphi_{j}\right\rangle&\left\langle\varphi_{\ell},\varphi_{k}\right\rangle&1\end{pmatrix}$
is constant for all $j\neq k\neq\ell\neq j$ is called $3$-uniform [BP05] (or $3_{c}$-uniform [HS12]). It was shown in [HS12] that the only $3_{c}$-uniform ETFs are, up to switching equivalence, orthonormal bases, simplices, or $(d,2d)$-ETFs with purely imaginary inner products. Thus, we would like to show that $3$-homogeneous ETFs are $3_{c}$-uniform. To that end, we compute the characteristic polynomial of the Gram submatrix above:
$\det\begin{pmatrix}1-\lambda&\left\langle\varphi_{j},\varphi_{k}\right\rangle&\left\langle\varphi_{j},\varphi_{\ell}\right\rangle\\ \left\langle\varphi_{k},\varphi_{j}\right\rangle&1-\lambda&\left\langle\varphi_{k},\varphi_{\ell}\right\rangle\\ \left\langle\varphi_{\ell},\varphi_{j}\right\rangle&\left\langle\varphi_{\ell},\varphi_{k}\right\rangle&1-\lambda\end{pmatrix}$
$=(1-\lambda)^{3}-(1-\lambda)\left(\left|\langle\varphi_{j},\varphi_{k}\rangle\right|^{2}+\left|\langle\varphi_{j},\varphi_{\ell}\rangle\right|^{2}+\left|\langle\varphi_{k},\varphi_{\ell}\rangle\right|^{2}\right)+\operatorname*{TP}(j,k,\ell)+\overline{\operatorname*{TP}(j,k,\ell)}$
$=(1-\lambda)^{3}-3(1-\lambda)\left|a\right|^{2}+2\operatorname{Re}a^{3},$
which is constant for any distinct $j,k,\ell$. Thus, the spectral norms of any of the $3\times 3$ Gram submatrices are equal to each other as their spectra are all equal. ∎
The concept of $k$-uniformity was introduced to study robustness of frames to $k$ erasures [HP04, BP05]. (Optimal total volume [CC23] of Parseval frames is also a related concept.) Equal-norm tight frames are optimal under certain models to noise and single erasures [GKK01] and are in turn $1$-uniform. ETFs are optimal to two erasures and are $2$-uniform [BP05, SH03, HP04]. ETFs with all zero, all negative, or all purely imaginary triple products are optimal to $3$ erasures and $3$-uniform [BP05, HS12]. Symmetry and $k$-uniformity are related, with $k$-fold symmetry almost implying $k$-uniformity, as outlined below. It is known that a Fourier frame, i.e., a frame formed as an all-column submatrix of a Fourier transform matrix is always an equal-norm tight frame with transitive vector symmetry group containing the group of modulations (similar to Proposition 2.8). More generally, the orbit of a non-zero vector under a finite irreducible unitary representation is always an equal norm tight frame [VW05] with transitive vector symmetry group. There are other results concerning when orbits of a non-zero vector under a projective unitary representation yields an equal-norm tight frame with necessarily transitive line symmetry group (see, e.g., [LPW05, ACGO25]). It is, however, possible for a spanning set to have a transitive vector or line symmetry group and not be a tight frame. For example, if one multiplies the first entry of every vector in a Paley ETF $\Phi_{p}$ by a non-unimodular scalar, the resulting set of vectors form a spanning set with a transitive vector/line symmetry group that contains cyclic permutations induced from the Paley modulation, but the new vectors would no longer be a tight frame / $1$-uniform frame. Any spanning set which has $2$-transitive line symmetry group is an ETF [IM24], and it immediately follows from the definition of $2$-homogeneity that any equal-norm tight frame which has $2$-homogeneous line symmetry group is an ETF, where ETFs are precisely the $2$-uniform frames. The proof of Lemma 3.3 shows that any $3$-homogeneous ETF must be $3$-uniform. ETFs which have purely imaginary triple products must have $n=2d$ [ET00, Str08] and always exist when $2d\times 2d$ skew-symmetric conference matrices exist [DGS91]. However, we will see in Theorem 3.5 that only certain skew-symmetric conference matrices will lead to $3$-homogeneous ETFs.
As we will see in Theorem 3.5, $k$-homogeneity of ETFs for $k>1$ only holds in a few interesting cases. For $\geq 4, kk$-homogeneity basically only holds for orthonormal bases and simplices or because (nk)(n-k)-homogeneity implies $k$-homogeneity. The remaining cases to study are $2$-homogeneous ETFs which are not $2$-transitive (otherwise one could look to [IM22, DK23, IM24]) and $2$-transitive ETFs which are $3$-homogeneous. The latter case is completely characterized below. The $(2,4)$-SIC is the unique (up to switching equivalence and vector permutation, cf. [Zhu10]) ETF of $4$ vectors in $\mathbb{C}^{2}$ and the real $(3,6)$-ETF is the unique (up to switching equivalence and vector permutation, cf. [BK06]) real $(3,6)$-ETF.
Theorem 3.5.
The only ETFs with $3$-homogeneous line symmetry groups are the $(2,4)$-SIC, the $(d,2d)$-ETFs with $d=q+1$ where $q\equiv 3\bmod 4$ is a prime power arising from Paley skew-conference matrices, orthonormal bases, and simplices, which are all $2$-transitive. The only ETFs with $4$-homogeneous line symmetry groups are the $(2,4)$-SIC, the real $(3,6)$-ETF, orthonormal bases, and simplices.